OpenStudy (nicole202):

What is 3,920,000,000,000 in scientific notation? 3.92×10103.92×1010 3.92×10123.92×1012 3.92×10−103.92×10−10 3.92×10−12

OpenStudy (rcp031):

3.92×10^(−12)

OpenStudy (rcp031):

scientific notation? means is a special way of writing numbers: How to Do it To figure out the power of 10, think "how many places do I move the decimal point?" When the number is 10 or greater, the decimal point has to move to the left, and the power of 10 is positive. When the number is smaller than 1, the decimal point has to move to the right, so the power of 10 is negative. Example: 0.0055 is written 5.5 × 10-3 Because 0.0055 = 5.5 × 0.001 = 5.5 × 10-3 Example: 3.2 is written 3.2 × 100 We didn't have to move the decimal point at all, so the power is 100 But it is now in Scientific Notation